# VLA-Pruner: Temporal-Aware Dual-Level Visual Token Pruning for Efficient Vision-Language-Action Inference

### v1

@@ A.2 Complexity Analysis
Following FastV [8], SparseVLM [60] and DivPrune [2], we estimate the computational requirement of VLA-Pruner in terms of Transformer FLOPs, and focus on how pruning visual tokens changes the effective sequence length seen by each layer. As in prior work [2], we count the FLOPs of multi-head attention and feed-forward networks (FFNs) in the backbone (the overhead of token selection as negligible compared to the main model FLOPs.) Let be the total number of Transformer layers in the VLA backbone and the index of the pruning layer. Denote the numbers of text and visual tokens before pruning by and (), and let be the full sequence length. After VLA-Pruner keeps a fraction of the visual tokens, the sequence length becomes . Following [8], the FLOPs of attention and FFN for one layer with sequence length , hidden size , and FFN intermediate size can be approximated as
@@ B.1.1 VLA Model Details
We consider three representative VLA architectures in our experiments: OpenVLA, OpenVLA-OFT, and . All models take RGB observations and a natural language instruction as input, and produce robot actions end-to-end. OpenVLA. OpenVLA [22] is a 7B-parameter open-source vision–language–action model built by fine-tuning a Prismatic VLM (Llama 2 language backbone with DINOv2 and SigLIP visual encoders) on robot manipulation episodes from the Open X-Embodiment dataset. The model represents actions as discrete tokens and decodes them autoregressively, enabling a single policy to control multiple embodiments out-of-the-box. OpenVLA serves as our base autoregressive VLA, and we follow its official checkpoints and evaluation settings on LIBERO. OpenVLA-OFT. OpenVLA-OFT [24] is an “Optimized Fine-Tuning” variant of OpenVLA that improves both control performance and inference efficiency. It augments the original decoder with parallel decoding and action chunking, and switches to a continuous action representation trained with a simple L1 regression objective. This substantially boosts success rates on LIBERO while increasing action-generation throughput. To be noted, OpenVLA-OFT utilizes bidirectional attention, meaning that during inference, there is already an action-to-vision cross-attention. However, we observe that during the early-layer prefill phase, the attention still focuses on semantic understanding, making action-to-vision less significant. In deeper layers, the action-to-vision attention starts to exhibit more distinct properties. This allows our method to effectively enhance token pruning performance by utilizing past timestep action-to-vision cross-attention at the later layers to guide current token pruning. Model. [4] is a prototype generalist robot policy that augments a pretrained vision–language model (with a PaliGemma backbone) with a continuous-action flow-matching head. Built on top of a large VLM to inherit Internet-scale semantic knowledge, is trained on diverse real-robot data and is designed for high-frequency control in dexterous manipulation tasks. The model decodes continuous motor commands via a flow-matching process with 10 integration steps. We use as a diffusion-based VLA backbone to evaluate the generality of VLA-Pruner. To capture the most informative action-to-vision interactions, we record the action-to-vision cross-attention averaged over the latter half of these integration steps, where the cross-attention becomes more stable and reliable.
@@ B.1.2 Acceleration Baseline Details
DivPrune. DivPrune [2](CVPR25) addresses the tendency of attention-based methods to select similar and redundant tokens. Instead of ranking tokens by attention scores, it formulates token selection as a Max–Min Diversity Problem and chooses a subset of visual tokens that are maximally dissimilar in feature space, thereby reducing redundancy while maintaining coverage. However, DivPrune is agnostic to the dual-level relevance of visual tokens in VLA prefilling and action decoding, and its diversity objective fall short for VLA tasks where localized details are more critical.
VLA-Cache. VLA-Cache [55](Neurips25) is a training-free acceleration method tailored for VLA models. It exploits temporal continuity in robotic manipulation by identifying visually static tokens whose features change little across timesteps, and reuses their computation via a KV-cache. At the same time, it uses text-to-vision cross-attention to keep task-relevant tokens fully computed. This cache-based mechanism reduces repeated computation over static background regions, but its token selection remains coarse-grained, which does not explicitly consider action-to-vision attention as we do. To ensure a fair comparison, we set the dynamic token limit based on similarity to be ( is the token budget), and use text-to-vision score to exclude of the most relevant tokens. These settings follow the default setup of VLA-Cache codebase.
EfficientVLA. EfficientVLA [56](Neurips25) proposes a structured acceleration framework for VLA models that combines visual token pruning, layer reduction, and diffusion temporal feature caching. Its token-pruning component leverages attention and diversity heuristics inherited from VLM pruning, while additional modules (i.e., layer skipping and diffusion caching) are used to achieve comparable acceleration. In contrast, VLA-Pruner is a single, plug-and-play token pruning module. These approaches are therefore complementary: EfficientVLA and other VLA acceleration methods can be adopted orthogonally with VLA-Pruner, while VLA-Pruner achieves strongest token-prune performance.
@@
| EfficientVLA (w/o layer) | 85.1 | 81.6 | 76.5 | 51.7 | 73.7 | 58.12% | 1.334 |
| EfficientVLA | 84.9 | 80.3 | 76.1 | 50.3 | 72.9 | 41.30% | 1.491 |
@@ B.2 Benchmark Details
SIMPLER [27] is a simulation suite built to closely track the real Google robot environment while allowing controlled appearance variation. It provides two evaluation settings. The Visual Matching (VM) setting keeps backgrounds, lighting, distractors, and camera poses close to the real setup so that rendered observations visually resemble real robot scenes and better predict real-world performance. The Variant Aggregation (VA) setting starts from the same base scenes but procedurally perturbs nuisance factors such as background textures, illumination, number and placement of distractor objects, table material, and camera viewpoints, creating a broad distribution shift for testing robustness. For the Google robot configuration, SIMPLER defines four canonical manipulation tasks: “pick coke can”,”pick object”, “move near”, “open/close Drawer”, and “place in closed drawer”. In our experiments, we follow the official simulator settings and success metrics, and report results on three representative tasks—Pick Coke Can, Move Near, and Open/Close Drawer—under VM or VA setting.
@@
Pruning layer . We compare results of different pruning layer values in Table 6. VLA-Pruner details. We show ratio trends at different pruning ratios within a robotic task in Fig. 10.
@@ B.6 More Visualization
We provide detailed visualization of VLA-Pruner token prune results across LIBERO tasks under diverse ratios in Fig. 11. VLA-Pruner maintains both local details for action execution and global information for semantic understanding and task planning.
@@
Our main contribution is to point out the dual-system nature of VLA inference and to propose a dual-level importance criterion that jointly accounts for semantic and action-level relevance. However, our current estimation of action-level importance relies on a heuristic temporal smoothing scheme with a fixed window and decay rate. While this works well for most tabletop settings, its benefits may diminish in highly dynamic scenarios (e.g., egocentric wrist-camera views), where rapid viewpoint shifts and object motion can violate the short-term temporal continuity assumption and blur sharp changes in action-to-vision attention. A promising direction for future work is to replace the fixed temporal smoothing with an adaptive prediction module that adjusts its effective window based on motion cues or feature changes (e.g., magnitude of visual dynamics or action-conditioned attention variance), or to learn a lightweight temporal attention network that directly predicts per-token importance from recent history. Such adaptive mechanisms may further improve pruning quality in dynamic environments while retaining the efficiency advantages of VLA-Pruner.

### v2

@@ A.2 Complexity Analysis
Following FastV [8], SparseVLM [60] and DivPrune [2], we estimate the computational requirement of VLA-Pruner in terms of Transformer FLOPs, and focus on how pruning visual tokens changes the effective sequence length seen by each layer. As in prior work [2], we count the FLOPs of multi-head attention and feed-forward networks (FFNs) in the backbone (the overhead of token selection as negligible compared to the main model FLOPs.) Let be the total number of Transformer layers in the VLA backbone and the index of the pruning layer (we prune tokens before the layer). Denote the numbers of text and visual tokens before pruning by and (), and let be the full sequence length. After VLA-Pruner keeps a fraction of the visual tokens, the sequence length becomes . Following [8], the FLOPs of attention and FFN for one layer with sequence length , hidden size , and FFN intermediate size can be approximated as
@@ B.1.1 VLA Model Details
We consider three representative VLA architectures in our experiments: OpenVLA, OpenVLA-OFT, and . All models take RGB observations and a natural language instruction as input, and produce robot actions end-to-end. OpenVLA. OpenVLA [22] is a 7B-parameter open-source vision–language–action model built by fine-tuning a Prismatic VLM (Llama 2 language backbone with DINOv2 and SigLIP visual encoders) on robot manipulation episodes from the Open X-Embodiment dataset. The model represents actions as discrete tokens and decodes them autoregressively, enabling a single policy to control multiple embodiments out-of-the-box. OpenVLA serves as our base autoregressive VLA, and we follow its official checkpoints and evaluation settings on LIBERO. OpenVLA-OFT. OpenVLA-OFT [24] is an “Optimized Fine-Tuning” variant of OpenVLA that improves both control performance and inference efficiency. It augments the original decoder with parallel decoding and action chunking, and switches to a continuous action representation trained with a simple L1 regression objective. This substantially boosts success rates on LIBERO while increasing action-generation throughput. To be noted, OpenVLA-OFT utilizes bidirectional attention, meaning that during inference, there is already an action-to-vision cross-attention. However, we observe that during the early-layer prefill phase, the attention still focuses on semantic understanding, making action-to-vision less significant. In deeper layers, the action-to-vision attention starts to exhibit more distinct properties. This allows our method to effectively enhance token pruning performance by utilizing past timestep action-to-vision cross-attention at the deeper layers to guide current token pruning. Model. [4] is a prototype generalist robot policy that augments a pretrained vision–language model (PaliGemma backbone) with a continuous-action flow-matching head. Built on top of a large VLM to inherit Internet-scale semantic knowledge, is trained on diverse real-robot data and is designed for high-frequency control in dexterous manipulation tasks. The model decodes continuous motor commands via a flow-matching process with 10 integration steps. We use as a diffusion-based VLA backbone to evaluate the cross-architecture generalizability of VLA-Pruner. To capture informative interactions and reduce noise, we record the action-to-vision cross-attention scores averaged over the latter half of these integration steps.
@@ B.1.2 Acceleration Baseline Details
DivPrune. DivPrune [2](CVPR25) addresses the tendency of attention-based methods to select similar and redundant tokens. Instead of ranking tokens by attention scores, it formulates token selection as a Max–Min Diversity Problem and chooses a subset of visual tokens that are maximally dissimilar in feature space, thereby reducing redundancy while maintaining coverage. However, DivPrune is agnostic to the dual-level relevance of visual tokens in VLA prefilling and action decoding, and its diversity objective falls short for VLA tasks where localized details are critical.
VLA-Cache. VLA-Cache [55](Neurips25) is a training-free acceleration method tailored for VLA models. It exploits temporal continuity in robotic manipulation by identifying visually static tokens whose features change little across timesteps, and reuses their computation via a KV-cache. At the same time, it uses text-to-vision cross-attention to keep task-relevant tokens fully computed. This cache-based mechanism reduces repeated computation over static background regions, but its semantic attention score based token selection remains coarse-grained, which overlooks distinct action attention distribution. To ensure a fair comparison, we set the dynamic token limit to be ( is the token budget), and use text-to-vision score to exclude of the most relevant tokens. These settings follow the default setup of VLA-Cache codebase.
EfficientVLA. EfficientVLA [56](Neurips25) proposes a structured acceleration framework for diffusion-based VLA models that combines visual token pruning, layer reduction, and diffusion temporal feature caching. Its token-pruning component leverages semantic attention and diversity heuristics inherited from VLM pruning, while additional modules (i.e., layer skipping and diffusion caching) are used to achieve comparable acceleration. In contrast, VLA-Pruner is a single, plug-and-play token pruning module. These approaches are therefore complementary: EfficientVLA and other VLA acceleration methods can be adopted orthogonally with VLA-Pruner, while VLA-Pruner achieves strongest token-prune performance.
@@
| EfficientVLA (T-only) | 84.3 | 81.6 | 76.5 | 49.8 | 73.1 | 58.12% | 1.334 |
| EfficientVLA | 83.2 | 80.3 | 75.7 | 49.1 | 72.0 | 41.30% | 1.491 |
@@ B.2 Benchmark Details
SIMPLER [27] is a simulation suite built to closely track the real Google robot environment while allowing controlled appearance variation. It provides two evaluation settings. The Visual Matching (VM) setting keeps backgrounds, lighting, distractors, and camera poses close to the real setup so that rendered observations visually resemble real robot scenes and better predict real-world performance. The Variant Aggregation (VA) setting starts from the same base scenes but procedurally perturbs nuisance factors such as background textures, illumination, number and placement of distractor objects, table material, and camera viewpoints, creating a broad distribution shift for testing robustness. For the Google robot configuration, SIMPLER defines four canonical manipulation tasks: “pick coke can”,“pick object”, “move near”, “open/close Drawer”, and “place in closed drawer”. In our experiments, we follow the official simulator settings and success metrics, and select three representative tasks—Pick Coke Can, Move Near, and Open/Close Drawer—under VM or VA setting, where base model attains non-trivial performance.
@@
Pruning layer . We compare results of different pruning layer values in Table 6 (retain 25% tokens). VLA-Pruner details. We show ratio trends at different pruning ratios for a robotic task in Fig. 10.
@@ B.6 More Visualization
We provide detailed visualization of VLA-Pruner token prune results across LIBERO tasks (50%, 75% pruning ratios) in Fig. 11. VLA-Pruner precisely maintains both local details for action execution and global information for semantic understanding and task planning.
@@
Our main contribution is to point out the dual-system nature of VLA inference and to propose a dual-level importance criterion that jointly accounts for semantic and action-level relevance. However, our current estimation of action-level importance relies on a heuristic temporal smoothing scheme with a fixed window and decay rate. While this works well for most tabletop settings, its benefits may diminish in highly dynamic scenarios (e.g., egocentric wrist-camera views or conveyor-belt environments), where rapid viewpoint shifts and object motion can violate the short-term temporal continuity assumption and blur sharp changes in action-to-vision attention. A promising direction for future work is to replace the fixed temporal smoothing with an adaptive prediction module that adjusts its effective window based on motion cues or feature changes (e.g., magnitude of visual dynamics or action-conditioned attention variance), or to learn a lightweight temporal attention network that directly predicts per-token importance from recent history. Such adaptive mechanisms may further improve pruning quality in dynamic environments while retaining the advantages of VLA-Pruner.